# Improving Generalization in Game Agents with Data Augmentation in Imitation Learning

### 21 Dec 2023 v2

@@ I Introduction
Current automated testing techniques rely primarily on model-based automated playtesting [1, 2], but recent developments in Reinforcement Learning (RL) [3, 4, 5] have demonstrated that it is possible to train game Artificial Intelligence (AI) agents to reach human-level performance in complex video games. These techniques can also train agents to explore game scenes and exploit environments to detect bugs and/or glitches in the geometry [6, 7], thereby improving automated playtesting.
@@
However, most techniques used for automated game testing suffer from an efficiency–generalization trade-off. Model-based and RL policies are known to easily overfit [8, 9], and achieving greater generalization requires numerous training samples. Imitation Learning (IL) was proposed to reduce the sample-inefficiency problem of RL [10, 8]. An IL agent learns from demonstrations rather than experience, allowing game designers to use prior knowledge to guide the agent towards its goal without additional knowledge of programming or machine learning. Even though IL can mitigate the sample-inefficiency problem, the generalization issue persists. To create an IL agent capable of solving situations that are outside of the distribution of the demonstration set, a large number of demonstration samples are necessary, which reduces efficiency.
@@
In this paper, we investigate how to improve the generalization of IL agents by reducing sample-inefficiency via data augmentation. While data augmentation is commonly used to address generalization and efficiency in computer vision problems [11, 12], recent applications of image-based [13] and feature-based [14] state spaces have shown promising results in both online and offline RL. Consequently, we propose a comprehensive study of the different data augmentation techniques in IL, with a focus on augmented feature-based state spaces, as game AI typically operates with feature values instead of images. We address the following research questions: Which augmentation techniques perform best in a game AI IL setting? Which consistently perform well in different use cases? And which are ineffective in a game AI IL setting? To answer these questions, we train agents with different augmentation combinations in a training environment and evaluate their generalization across four test environments. This study shows that certain combinations are more suitable for this setting than others, with the best achieving a performance $1.6$ times higher than that of the non-augmented agent.
@@
This study aims to provide guidance to those seeking to use data augmentation to increase generalization in IL agents, particularly in the game testing context. It can also be used as a potential starting point for improving generalization of any feature-based game AI agent.
@@ II Related Work
While generalization has been widely discussed in reinforcement learning literature, only a few works have focused on generalization and imitation learning. In this section, we review the most relevant literature related to the present work.
@@ II-A Imitation Learning and Games
IL has emerged as a widely accepted method for creating agents to perform tasks in 3D video games without RL [8]. Amiranashvili et al. [18] used an agent trained solely through BC to play the game Minecraft. Chang et al. [19] used demonstrations to guide exploration, while Zhao et al. [20] employed a BC approach to teach agents based on game designers’ expertise. Harmer et al. [21] trained agents using a combination of IL and RL with multi-action policies for a first person shooter game. Additionally, Tucker et al. [22] used an inverse RL technique to train agents capable of playing a variety of games in the Arcade Learning Environment suite [23]. Other notable examples include the work of Sestini et al. [10], who combined RL, IL, and curiosity-driven learning to train playtesting agents, and Ferguson et al. [24], who proposed the use of dynamic time warping IL to imitate distinct playstyles. However, all of these models must address the efficiency–generalization trade-off: higher levels of generalization require more demonstrations or more interactions with the environment. In this paper we examine the latest data augmentation techniques used in RL to increase generalization and efficiency and apply them in IL.
@@ II-B Generalization in Reinforcement Learning
In several use cases, policies produced with RL, IL or inverse reinforcement learning have overfit to their training environments [25, 26]. Techniques like domain randomization and procedural content generation improve generalization and help prevent overfitting [27, 28]. In inverse RL, Sestini et al. [26] proved that with the correct training setup, domain randomization can mitigate overfitting and improve efficiency. However, domain randomization requires having a procedurally ready game or access to the environment code to implement randomized elements. In practical settings such as video game production, domain randomization is not always feasible.
Data augmentation has also been investigated in the context of RL and generalization [29]. Laskin et al. [30] proposed a thorough survey regarding data augmentation and RL for image data, while Zolna et al. [31] combined data augmentation and the GAIL algorithm to train a sample-efficient IL agent in a robotic setting. The present work takes inspiration from the S4RL algorithm of Sinha et al. [14]. These authors investigated the efficacy of using data augmentation in offline reinforcement learning on a feature-based state space. They proposed $7$ different augmentation schemes and analyzed their performance with existing offline reinforcement learning algorithms. Furthermore, they combined the most successful data augmentation scheme with a state-of-the-art Q-learning technique. Similarly, this present work investigates the effects of $6$ feature-based data augmentation techniques for training game-playing agents with IL.
@@ III Method
In this section we first review the main IL technique, behavioral cloning, and discuss the rationale behind its selection. After that, we review data augmentations previously studied in offline RL research and explain how we adapt these techniques to improve sample efficiency and generalization in our IL agent use case.
@@ III-A Behavioral Cloning
We begin by introducing the main algorithm used in this study. Since the focus is on data augmentation rather than selecting the optimal algorithm, we choose one of the simplest IL algorithms, BC. Recent literature has shown BC to be a good algorithm for IL in 3D games [8]. Formally, we describe our domain as a Markov Decision Process (MDP) consisting of a state space $S\in\mathbb{R}^{n}$, an action space $A\in\mathbb{R}^{m}$, a transition dynamic function $p(s^{\prime}|s,a):S\times A\rightarrow p(S)$ and a parameterized policy $\pi_{\theta}:S\rightarrow p(A)$ that is a mapping from the state space $S$ to a probability distribution over the set of actions $A$. Section IV provides a complete description of the particular state- and action-space used in this work. Note that in our setting we do not have a reward function since the agent is learning from expert demonstrations. Given a set of optimal demonstrations $\mathcal{D}=\{\tau_{i}\;|\;\tau_{i}=(s_{0}^{i},a_{0}^{i},...,s_{T_{i}}^{i},a_{T_{i}}^{i}),\;i=1,..,N\}$ of $N$ trajectories $\tau_{i}$, each composed of a sequence of state-action pairs $(s^{i}_{k},a^{i}_{k})$, we define the BC objective used to update the network as:
@@
This objective aims to mimic the expert behavior, which is represented by the dataset $\mathcal{D}$.
@@ IV-A Environment and State Space
The environment used in this study is an open-world city simulation, as illustrated in Figure 1(a) and Figure 1(b). The agent has a discrete action space of size $9$, consisting of moving forward, moving backward, right rotation, left rotation, jumping, shooting, right sidestep, left sidestep, and no action. The state space was proposed by Sestini et al. [8] and consists of a game goal position, represented as the $\mathbb{R}^{2}$ projections of the agent-to-goal vector onto the $XY$ and $XZ$ planes, normalized by the gameplay area size, along with game-specific state observations such as the agent’s climbing status, contact with the ground, presence in an elevator, jump cooldown, and weapon magazine status. Observations includes a list of entities and game objects that the agent should be aware of, e.g., intermediate goals, dynamic objects, enemies, and other assets that could be useful for achieving the final goal. For these entities, the same relative information from agent-to-goal is referenced, except as agent-to-entity. For our use case, the entities are the button to open the door and the goal itself. Additionally, a 3D semantic map is used for local perception. This map is a categorical discretization of the space and elements around the agent. Each voxel in the map carries a semantic integer value describing the type of object at the corresponding game world position. For this work, we use a semantic map of size $5\times 5\times 5$. In this environment, an episode consists of a maximum of $750$ steps. An episode is marked a success if the agent reaches the goal before the timeout. As we discuss in Section V-B, from this game we create a training environment and a testing environment.
@@ IV-C Training Setup
Training is a supervised learning problem where the input data for each model is the unique dataset described in Section V-A and the output is an agent’s action as described in Section IV-A. We use a neural network, as described in Section IV-B, built with the TensorFlow framework. Each model is trained for $300$ epochs on a single machine with an AMD Ryzen Threadripper 1950X 16-Core Processor and an NVIDIA GV100 GPU, with an average training time of $336$ seconds.
@@ V Experiments and Results
In the training environment, we began by creating a set of demonstrations presenting the agent with the desired goal: the agent must reach a building, press a button to open a door, and enter the building before the door closes in order to reach the goal position. The complexity was relatively low, with few obstacles between the agent’s starting position and the goal. For the human player, there were two natural paths to reach the goal: diagonally through the park, or down the road and to the left. We used both paths to increase the range of possible states that the agent could encounter. We alternated between these two paths to maintain a balanced dataset for supervised learning. We provided human demonstrations totaling $78$ episodes, or $15,380$ samples.
For the test environments, we created $4$ distinct new environments from the initial training environment by making design changes to the original. This included moving and rotating the goal building, increasing the number of obstacles, and obstructing one of the direct paths to the goal building. For each environment, the agent had to complete the same task of navigating to the building, opening the door, and entering the goal position. A summary of design changes and subjective complexity is available in Table I and a visual for the modified environment is presented in Figure 1(c).
@@ V-A Augmented Models
As a baseline, we used the dataset consisting of the original $78$ episodes of data without any applied augmentations. Using this data and the augmentations, we created different datasets, each featuring one or more augmentations. We refer to the models trained using these datasets as augmented models, and they are labeled with the corresponding dataset’s augmentations. We applied all the augmentations listed in Section III-B and all their combinations, up to $3$ augmentations, i.e, one dataset could be augmented with $1$ augmentation, $2$ augmentations, or $3$ augmentations. To avoid having too many augmentation scenarios with only noise, we excluded the datasets that have both Gaussian and uniform noise. This process resulted in a total of $38$ different combinations of data augmentations.
@@
Figure 2: Overview of the augmentation process. We started with the original set of demonstrations. Then, depending on the number and type of the selected augmentations, we created a new, unique dataset. For example, if we selected two augmentations (Gaussian noise and semantic dropout), we started from the original dataset, applied Gaussian noise to it, and then augmented the resulting dataset with semantic dropout.
Dataset size and standard deviation of Gaussian noise are two significant parameters of augmentation. To better understand their effects, we conduct a hyperparameter study over these parameters. All hyperparameters together with the variations are summarized in Table II. With these variations, we trained a total of $228$ different models. For each model, we repeated the experiments for $10$ different seeds.
TABLE II: Summary of available hyperparameters. Values in square brackets were experimentally varied; all others remained fixed.
@@
V-B Evaluation
During testing, agents run in each environment for $100$ episodes. Episodes are reset after a timeout of $750$ steps. The episode is recorded as a success if the agent reaches the goal before the timeout. For an initial evaluation, we run the full set of $228$ models in the initial training environment. Using the mean success rate over $10$ seeds of each model as a selection criterion, we narrow the set of models down to $40$ to evaluate generalization performance in the testing environments. This set of models consists of the following two groups: 1) the $19$ top-performing models and 2) a selection of $20$ of the lowest-performing models. For the lowest-performing group, we selected models with relative success rates ranging from $0.00$ to $0.39$ to obtain a mix of both low-performing and failing models. In addition, we selected the highest-performing baseline model and included it in the top-performing group. Our main focus is to evaluate how data augmentation can impact the generalization performance of the trained agents. For this reason, we evaluate only the top $19$ models found in the initial evaluation. We are interested in finding the relative success rate of each augmented model compared to the baseline model for each test environment. This provides a quantitative measure of the improvements, if any, that are achieved through the data augmentation combination, so we use this as our primary benchmark. We also seek to identify consistency in model performance across environments, as it might suggest a clear set of combinations to use—or to avoid—in similar scenarios, as a starting point for other game environments.
Our first assessment is whether data augmentations could improve an imitation learning agent’s ability to generalize to unseen changes in an environment. The results, summarized in Table III, show that these augmented models are able to outperform the baseline in each of the test environments. Note that a mean relative success rate of 1.X corresponds to an improvement of X%. Hence, the mean relative success rates showed an improvement of $20\%$ up to $80\%$ for different environments but with a relatively high variance over different models.
@@ V-B Evaluation
The next assessment is to determine whether a combination of augmentations consistently improves generalization regardless of the environment. Figure 4 shows that $5$ different models outperform their respective baselines in $4$ of the $5$ environments. These $6$ models each includes a combination of at least $2$ augmentations and use at least $80\%$ of the data. While these models are the most consistent in outperforming the baseline, they do not have the $5$ highest relative success rates, which suggests that there is a trade-off between best achievable generalization performance and consistency over all testing environments.
Figure 4: Models that consistently outperform the baseline in the $4$ different testing environments plus the training environment. For the abbreviations legend, see Figure 3. Additional details about the augmentations are provided in Section III.

### 07 Apr 2024 v3

@@ I Introduction
Current automated testing techniques rely primarily on model-based automated playtesting [1, 2], but recent developments in Reinforcement Learning (RL) [3, 4, 5] have demonstrated that it is possible to train game Artificial Intelligence (AI) agents to reach human-level performance in complex video games, requiring however a vast amount of resources to achieve such training. These techniques can also train agents to explore game scenes and exploit environments to detect bugs [6, 7], thereby improving automated playtesting.
@@
However, most techniques used for automated game testing suffer from an efficiency–generalization trade-off. Model-based and RL policies are known to easily overfit [8, 9], and achieving greater generalization requires numerous training samples. Imitation Learning (IL) was proposed to reduce the sample-inefficiency problem of RL [10, 8]. An IL agent learns from demonstrations rather than experience, allowing game designers to use prior knowledge to guide the agent towards its goal without additional knowledge of programming or machine learning. Even though IL can mitigate the sample-inefficiency problem, the generalization issue persists. To develop an IL agent that can address situations beyond the demonstration set and minimize the distributional shift issue—-the phenomenon where the distribution of states and actions encountered by the learning agent differs from those in the expert demonstration data, leading to sub-optimal performance—a significant number of datasamples are needed. However, producing a large demonstration set can be costly. Additionally, even with a considerable amount of data, IL might still overfit to the training environment. Ideally, we should test the IL policy across various versions of the game, simulating game developers modifying it. Incorporating priors for new environments may be challenging, as developers may not anticipate future changes. Hence, it is crucial to have a model that can generalize beyond the training environment.
@@
In this paper, we investigate how to improve the generalization of IL agents by reducing sample-inefficiency via data augmentation. While data augmentation is commonly used to address generalization and efficiency in computer vision problems [11, 12], recent applications of image-based [13] and feature-based [14] state spaces have shown promising results in both online and offline RL. Consequently, we propose a comprehensive study of the different data augmentation techniques in IL, with a focus on augmented feature-based state spaces, as game AI typically operates with feature values instead of images. We address the following research questions: Which augmentation techniques perform best in a game AI IL setting? Which consistently perform well in different use cases? And which are ineffective in a game AI IL setting? To answer these questions, we train agents with different augmentation combinations in a training environment and evaluate their generalization across four test environments. We do not propose any new data augmentation techniques; instead, we conduct a comprehensive study of existing ones used in various domains and present an analysis of the combination of augmentations that yield the best results. This study shows that certain combinations are more suitable for this setting than others, with the best achieving a performance $1.6$ times higher than that of the non-augmented agent.
@@
Our main contribution aims to provide guidance to those seeking to use data augmentation to increase generalization in IL agents, particularly in the game testing context. It can also be used as a potential starting point for improving generalization of any feature-based game AI agent.
@@ II Related Work
While generalization has been widely discussed in reinforcement learning literature [9, 15], only a few works have focused on generalization and imitation learning. In this section, we review the most relevant literature related to this work.
@@ II-A Imitation Learning and Games
IL has emerged as a widely accepted method for creating agents to perform tasks in 3D video games without RL [8, 19]. Amiranashvili et al. [20] used an agent trained solely through BC to play the game Minecraft. Chang et al. [21] used demonstrations to guide exploration, while Zhao et al. [22] employed a BC approach to teach agents based on game designers’ expertise. Harmer et al. [23] trained agents using a combination of IL and RL with multi-action policies for a first person shooter game. Additionally, Tucker et al. [24] used an inverse RL technique to train agents capable of playing a variety of games in the Arcade Learning Environment suite [25]. Other notable examples include: the work of Sestini et al. [10], who combined RL, IL, and curiosity-driven learning to train playtesting agents; the work of Ferguson et al. [26], who proposed the use of dynamic time warping IL to imitate distinct playstyles; and the work of Pearce et al. [27], who introduce several innovations to make diffusion models suitable for imitating human behaviors in a first-person shooter game. However, all of these models must address the efficiency–generalization trade-off: higher levels of generalization require more demonstrations or more interactions with the environment. In this paper we examine the latest data augmentation techniques used in RL to increase generalization and efficiency and apply them in IL.
@@ II-B Generalization in Reinforcement Learning
In several use cases, policies produced with RL, IL or inverse RL have overfit to their training environments [15, 28]. Techniques like domain randomization and procedural content generation improve generalization and help prevent overfitting [29, 30]. In inverse RL, Sestini et al. [28] proved that with the correct training setup, domain randomization can mitigate overfitting and improve efficiency. However, domain randomization requires having a procedurally ready game or access to the environment code to implement randomized elements. In practical settings such as video game production, domain randomization is not always feasible.
Data augmentation has also been investigated in the context of RL and generalization [31]. Laskin et al. [32] proposed a thorough survey regarding data augmentation and RL for image data, while Zolna et al. [33] combined data augmentation and the GAIL algorithm to train a sample-efficient IL agent in a robotic setting. The present work takes inspiration from the S4RL algorithm of Sinha et al. [14]. These authors investigated the efficacy of using data augmentation in offline reinforcement learning on a feature-based state space. They proposed $7$ different augmentation schemes and analyzed their performance with existing offline reinforcement learning algorithms. Furthermore, they combined the most successful data augmentation scheme with a state-of-the-art offline RL technique. Similarly, this present work investigates the effects of $6$ feature-based data augmentation techniques for training game-playing agents with IL.
@@ III Method
Here we first review the main IL technique, behavioral cloning, and discuss the rationale behind its selection. Second, we review data augmentations previously studied in offline RL research and explain how we adapt these techniques to improve sample efficiency and generalization in our use case.
@@ III-A Behavioral Cloning
We begin by introducing the main algorithm used in this study. Since the focus is on data augmentation rather than selecting the optimal algorithm, we choose one of the simplest IL algorithms, BC. Recent literature has shown BC to be a good algorithm for IL in 3D games [8, 19]. Formally, we describe our domain as a Markov Decision Process (MDP) consisting of a state space $S\in\mathbb{R}^{n}$, an action space $A\in\mathbb{R}^{m}$, a transition dynamic function $p(s^{\prime}|s,a):S\times A\rightarrow p(S)$ and a parameterized policy $\pi_{\theta}:S\rightarrow p(A)$ that is a mapping from the state space $S$ to a probability distribution over the set of actions $A$. Section IV provides a complete description of the particular state- and action-space used in this work. Note that in our setting we do not have a reward function since the agent is learning from expert demonstrations. Given a set of optimal demonstrations $\mathcal{D}=\{\tau_{i}\;|\;\tau_{i}=(s_{0}^{i},a_{0}^{i},...,s_{T_{i}}^{i},a_{T_{i}}^{i}),\;i=1,..,N\}$ of $N$ trajectories $\tau_{i}$, each composed of a sequence of state-action pairs $(s^{i}_{k},a^{i}_{k})$, we define the BC objective used to update the network as:
@@
This equation follows the maximum entropy objective: by increasing the log-probability of the policy $\pi(a_{t}|s_{t})$ for a given $(a_{t},s_{t})\sim D$ pair sampled from the expert dataset $\mathcal{D}$, we enhance the probability of sampling that action in that specific state or similar states in proximity to $s_{t}$. At optimality, the policy $\pi$ mimics the expert behavior represented by $\mathcal{D}$.
@@ IV-A Environment and State Space
The environment used in this study is the same as used by Sestini et al. [8]. The game is an open-world city simulation, as illustrated in Figure 1(a) and Figure 1(b). The agent has a discrete action space of size $9$, consisting of moving forward, moving backward, right rotation, left rotation, jumping, shooting, right sidestep, left sidestep, and no action. In our experiments, as we eliminate the presence of enemies from the environment, we do not utilize the shooting action. The state space is the same proposed by Sestini et al. [8] and consists of a goal position, represented as the $\mathbb{R}^{2}$ projections of the agent-to-goal vector onto the $XY$ and $XZ$ planes, normalized by the gameplay area size, along with game-specific state observations such as the agent’s climbing status, contact with the ground, presence in an elevator, jump cooldown, and weapon magazine status. Observations includes a list of entities and game objects that the agent should be aware of, e.g., intermediate goals, dynamic objects, enemies, and other assets that could be useful for achieving the final goal. For these entities, the same relative information from agent-to-goal is referenced, except as agent-to-entity. For our use case, the entities are the button to open the door and the goal itself. Plus, a 3D semantic map is used for local perception. This map is a categorical discretization of the space around the agent. Each voxel in the map carries a semantic integer value describing the type of object at the corresponding game world position. We use a semantic map of size $5\times 5\times 5$. In this environment, an episode consists of a maximum of $750$ steps. An episode is marked a success if the agent reaches the goal before the timeout. As we discuss in Section V-B, from this game we create a training environment and different testing ones.
@@ IV-C Training Setup
Training is a supervised learning problem where the input data for each model is the unique dataset described in Section V-A. We use a neural network, as described in Section IV-B, built with the TensorFlow framework. Each model is trained for $300$ epochs on a single machine with an AMD Ryzen Threadripper 1950X 16-Core Processor and an NVIDIA GV100 GPU, with an average training time of $336$ seconds.
@@ V Experiments and Results
In the training environment, we began by creating a set of demonstrations presenting the agent with the desired goal: the agent must reach a building, press a button to open a door, and enter the building before the door closes in order to reach the goal position. This task is the same Use Case 1 in the work by Sestini et al. [8]. The complexity was relatively low, with few obstacles between the agent’s starting position and the goal. For the human player, there were two natural paths to reach the goal: diagonally through the park, or down the road and to the left. We used both paths to increase the range of possible states that the agent could encounter. We alternated between these two paths to maintain a balanced dataset for supervised learning. We provided human demonstrations totaling $78$ episodes, or $15,380$ samples.
For the test environments, we created four distinct new environments from the initial training environment by making design changes to the original. This included moving and rotating the goal building, increasing the number of obstacles, and obstructing one of the direct paths to the goal building. For each environment, the agent had to complete the same task of navigating to the building, opening the door, and entering the goal position. Herein, we describe the list of modifications for each test environment:
@@
Test Environment 1  - the position of the building containing the goal is slightly different from the training environment, with a slight rotation to the left. Additionally, we introduce $5$ more obstacles in the path from the agent’s starting position to the goal. We categorize the subjective difficulty of this variation as easy, consistent with the training environment;
Test Environment 2  - The position of the building containing the goal differs slightly along the x-axis, while the rotation remains the same as in the training environment. However, a total of $17$ obstacles obstruct the path from the agent’s starting position to the goal. We categorize the subjective difficulty of this variation as medium;
Test Environment 3  - The position of the building is significantly different from the one in the training environment, as can be seen by comparing Figure 1(a) and (c). Furthermore, the building has a different rotation, with a higher degree compared to Test Environment 1. This environment also includes a total of $5$ obstacles. We categorize the subjective difficulty of this variation as medium; and
Test Environment 4  - The position of the building is entirely different from the training environment. In contrast to the training environment, where the goal is in front of the agent facing its direction, the goal in this test environment is behind the agent’s starting position, facing its back. This configuration requires the agent to follow a completely different trajectory to reach the goal, one that is not in the expert dataset. Moreover, there are a total of $5$ obstacles. Due to the significant difference in the goal position, we categorize the subjective difficulty of this variation as hard.
A summary of design changes and subjective complexity is available in Table I and a visual for the modified environment is presented in Figure 1(c).
@@ V-A Augmented Models
As a baseline, we used the dataset consisting of the original $78$ episodes of data without any applied augmentations. Using this data and the augmentations, we created different datasets, each featuring one or more augmentations. We refer to the models trained using these datasets as augmented models, and they are labeled with the corresponding dataset’s augmentations. We applied all the augmentations listed in Section III-B and all their combinations, up to $3$ augmentations, i.e, one dataset could be augmented with $1$ augmentation, $2$ augmentations, or $3$ augmentations. Preliminary experiments showed that combining Gaussian and uniform noise together decreased the overall performances. For this reason, we excluded the datasets that have both Gaussian and uniform noise. This process resulted in a total of $38$ different combinations of data augmentations.
@@
Figure 2: Overview of the augmentation process. We start with the original set of demonstrations. Then, depending on the number and type of the selected augmentations, we create a new, unique dataset. For example, if we select two augmentations (Gaussian noise and semantic dropout), we start from the original dataset, apply Gaussian noise to it, and then augment the resulting dataset with semantic dropout.
Dataset size and standard deviation of Gaussian noise are two significant parameters. To better understand their effects, we conduct a hyperparameter study over these parameters. The tested ranges were selected following preliminary experiments. All hyperparameters together with the variations are summarized in Table II. With these variations, we trained a total of $228$ different models. Each model is a different combinations of augmentations, varying their hyperparameters. For instance, one model could be applying Gaussian noise with $\theta=0.03$ plus state mixup to $50\%$ of the original data. As we previously said, we have a total of $38$ combinations times the $6$ different sizes of the dataset. For each model, we repeated the experiments for $10$ different seeds.
TABLE II: Summary of available hyperparameters. Values in square brackets were experimentally varied; all others remained fixed. The ranges were selected following preliminary experiments.
@@
Figure 4: Models that consistently outperform the baseline in the $4$ different testing environments plus the training environment. For the abbreviations legend, see Figure 3.
V-B Evaluation
During testing, agents run in each environment for $100$ episodes. Episodes are reset after a timeout of $750$ steps. The episode is recorded as a success if the agent reaches the goal before the timeout. For an initial evaluation, we run the full set of $228$ models in the initial training environment. Using the mean success rate over $10$ seeds of each model as a selection criterion, we narrow the set of models down to $40$ to evaluate generalization performance in the testing environments. This set of models consists of the following two groups: 1) the $19$ top-performing models and 2) a selection of $20$ of the lowest-performing models. For the lowest-performing group, we selected models with relative success rates ranging from $0.00$ to $0.39$ to obtain a mix of both low-performing and failing models. In addition, we selected the highest-performing baseline model and included it in the top-performing group. Our main focus is to evaluate how data augmentation can impact the generalization performance of the trained agents. For this reason, we evaluate only the top $19$ models found in the initial evaluation. We are interested in finding the relative success rate of each augmented model compared to the baseline for each test environment. This provides a quantitative measure of the improvements, if any, that are achieved through the data augmentation combination, so we use this as our primary benchmark. We also seek to identify consistency in model performance across environments, as it might suggest a clear set of combinations to use—or to avoid—in similar scenarios, as a starting point for other game environments.
Our first assessment is whether data augmentations could improve an imitation learning agent’s ability to generalize to unseen changes in an environment. The results, summarized in Table III, show that these augmented models outperform the baseline in each of the test environments. Note that a mean relative success rate of 1.X corresponds to an improvement of X%. Hence, the mean relative success rates showed an improvement of $20\%$ up to $80\%$ for different environments but with a relatively high variance over different models.
The next assessment is to determine whether a combination of augmentations consistently improves generalization regardless of the environment. Figure 4 shows that $6$ different models outperform their respective baselines in $4$ of the $5$ environments. These $6$ models each includes a combination of at least $2$ augmentations and use at least $80\%$ of the data. While these models are the most consistent in outperforming the baseline, none of them have the highest relative success rates, but $3$ of them are in the top $6$ in the same evaluation (see Figure 3). This suggests that there is a trade-off between best achievable generalization performance and consistency over all testing environments.